# Finite-Agent Stochastic Differential Games on Large Graphs: I. The Linear-Quadratic Case

### v1

@@ 5.3 The time complexity analysis and running time comparison
For illustration purpose, the analysis of time complexity uses the forward Euler method as the numerical ODE solver. In the baseline construction for solving the Riccati system (2.13), each Euler step computes $(\check{F}^{1}_{t+h},\ldots,\check{F}^{N}_{t+h})$ based on $(\check{F}^{1}_{t},\ldots,\check{F}^{N}_{t})$. The calculation of the derivative $\dot{\check{F}}^{i}_{t}$ requires $O(N^{3})$ calculations, so a single Euler step costs $O(N^{4})$, and constructing the entire baseline solution requires $O(N_{T}N^{4})$. Meanwhile, only one matrix-valued ODE is numerically solved during the semi-explicit construction, so the whole procedure only costs $O(N_{T}N^{3})$.
@@
[…]
@@ 6 Conclusion and Future Studies
As machine learning methods have become more powerful, there has been an increasing trend to solve control problems and games numerically using deep learning and reinforcement learning techniques. We refer readers to the survey [HL23] for a general overview of these methods. In particular, direct parameterization [HE16] and Deep BSDE [EHJ17] are two popular methods for control problems and have been generalized to various extents. Furthermore, [HH20, Hu21] use the deep version of fictitious play as a fundamental strategy for solving the NE. Along with this research direction, we are developing graph-based deep-learning algorithms for solving the Nash equilibrium of large sparse network games. This effort is mainly motivated by the theoretical results and numerical findings in this work. The algorithm accelerates training by reducing the number of parameters without significantly losing the accuracy. The details of the algorithm and the numerical outcomes will be presented in an upcoming work [HLZ].
Finally, we point out that several directions are worth further investigation. For instance, although the global convergence of fictitious play has been observed in numerical experiments even when all model parameters take relatively large values, we were only able to provide the proof of convergence (cf. Corollary 3.8) under a smallness condition. Further relaxation of the technical assumption would be valuable. Additionally, regarding the rate of convergence, fictitious play numerically exhibits a graph-dependent rate as seen in Section 5.2, which has not been accurately characterized in Corollary 3.8. These studies on the convergence of fictitious play are left to future investigation. Separately, for linear-quadratic games on vertex-transitive graphs, we encounter technical difficulties when developing the semi-explicit characterization for the NE when $\varepsilon\neq q^{2}$, and it remains unsolved. The game equilibrium and asymptotic correlations of the limiting regime $(N\to\infty)$, in the spirit of [LS22], are also interesting topics for future work.
@@ Acknowledgement
R.H. was partially supported by a grant from the Simons Foundation (MP-TSM-00002783).

### v2

@@ 5.3 The time complexity analysis and running time comparison
For illustration purposes, the analysis of time complexity uses the forward Euler method as the numerical ODE solver. In the baseline construction for solving the Riccati system (2.14), each Euler step computes $(\check{F}^{1}_{t+h},\ldots,\check{F}^{N}_{t+h})$ based on $(\check{F}^{1}_{t},\ldots,\check{F}^{N}_{t})$. The calculation of the derivative $\dot{\check{F}}^{i}_{t}$ requires $O(N^{3})$ calculations, so a single Euler step costs $O(N^{4})$, and constructing the entire baseline solution requires $O(N_{T}N^{4})$. Meanwhile, only one matrix-valued ODE is numerically solved during the semi-explicit construction, so the whole procedure only costs $O(N_{T}N^{3})$.
@@
where the initial state $x_{0}$ is randomly sampled from a uniform distribution $x_{0}^{i}\overset{\text{i.i.d.}}{\sim}\mathcal{U}(-1,1),\ \forall i\in[N]$. The remaining model parameter $N$ varies and will be specified in Table 5. To ensure a fair comparison, we fix the error tolerance level at $\text{tol}=10^{-5}$ as indicated by Table 1, and stop the FP procedure at stage $k$ once $\text{MAE}(\check{F}^{\cdot,k-1},\check{F}^{\cdot,k})<\text{tol}$. The numerical outcomes are presented in Table 5 and align with the time complexity analysis shown in Table 4.
@@ 6 Conclusion and Future Studies
As machine learning methods have become more powerful, there has been an increasing trend to solve control problems and games numerically using deep learning and reinforcement learning techniques. We refer readers to the survey [HL23] for a general overview of these methods. In particular, direct parameterization [HE16] and Deep BSDE [EHJ17] are two popular methods for control problems and have been generalized to various extents. Furthermore, [HH20, Hu21] use the deep version of fictitious play as a fundamental strategy for solving the NE. Along with this research direction, we are developing graph-based deep-learning algorithms for solving the Nash equilibrium of large sparse network games. This effort is mainly motivated by the theoretical results and numerical findings in this work. The algorithm accelerates training by reducing the number of parameters without significantly losing the accuracy. The details of the algorithm and the numerical outcomes will be presented in our upcoming work [HLZ].
Finally, we point out that several directions are worth further investigation. For instance, although the global convergence of fictitious play has been observed in numerical experiments even when all model parameters take relatively large values, we were only able to provide the proof of convergence (cf. Corollary 3.8) under a smallness condition. Further relaxation of the technical assumption would be valuable. Additionally, regarding the rate of convergence, fictitious play numerically exhibits a graph-dependent rate as seen in Section 5.2, which has not been accurately characterized in Corollary 3.8. These studies on the convergence of fictitious play are left to future investigation. Separately, for linear-quadratic games on vertex-transitive graphs, we encounter technical difficulties when developing the semi-explicit characterization for the NE when $\varepsilon\neq q^{2}$, which remains unsolved. The game equilibrium and asymptotic correlations of the limiting regime $(N\to\infty)$, in the spirit of [LS22], are also interesting topics for future work. In addition, analyzing our generic model under more restrictive information structures, e.g., a partially observable setting where agents only access local information within its graph neighborhood and maintain beliefs for unobservable information, would be an interesting direction for future work. Along this research direction, numerical works have been done in the field of multi-agent reinforcement learning under a collaborative game setting [CCK20, ZYL+18]. Theoretical works have been done under a collaborative setting in the mean-field regime when a network of states (but not players) is given [GGWX25]. To our best knowledge, there is no theoretical analysis for the setting of a competitive game on graphs with local information structures. Such a setting provides a more realistic modeling of strategic interactions on graphs, compared to the complete information case discussed in this paper, and deserves future investigations.
@@ Acknowledgement
R.H. was partially supported by a grant from the Simons Foundation (MP-TSM-00002783), and the ONR grant under #N00014-24-1-2432.
Appendix A Heuristic Derivations
This appendix provides detailed heuristic derivations of how equations (4.5), together with (4.3)–(4.4) and (4.6), are obtained. This supplements the steps briefly outlined in Section 4.1.2.
During the heuristic derivations below, several assumptions must be made. For a smoother presentation flow, these assumptions (A.1)–(A.4) are summarized later in Remark A.6, and are referred to during the heuristic derivation. However, we emphasize that verifying these assumptions is not necessary due to the existence of the verification procedure presented in Section 4.3.
Let $P:[0,T]\to\mathbb{R}^{N\times N}$ be a matrix-valued function that is defined in terms of $F^{1}_{t},\cdots,F^{N}_{t}$:
$P_{t}:=\sum_{k=1}^{N}\left[(a+q)L+F^{k}_{t}\right]e_{k}e_{k}^{\operatorname{T}}=(a+q)L+\sum_{k=1}^{N}F^{k}_{t}e_{k}e_{k}^{\operatorname{T}}.$ (A.1)
$\dot{F}^{i}_{t}-P_{t}F^{i}_{t}-F^{i}_{t}P_{t}+(\varepsilon-q^{2})Le_{i}e_{i}^{\operatorname{T}}L+F^{i}_{t}e_{i}e_{i}^{\operatorname{T}}F^{i}_{t}=0,\quad F^{i}_{T}=cLe_{i}e_{i}^{\operatorname{T}}L.$ (A.2)
Treating $P$ as given, the solution to equation (A.2) can be represented in terms of $P_{t}$, denoted as $F^{1}_{t}(P),\cdots,F^{N}_{t}(P)$. Substituting $F^{1}_{t}(P),\cdots,F^{N}_{t}(P)$ into equation (A.1) provides an equation solely for $P$:
$P_{t}=(a+q)L+\sum_{k=1}^{N}F^{k}_{t}(P)e_{k}e_{k}^{\operatorname{T}}.$ (A.3)
Definition A.1.
$R_{\varphi}e_{i}=e_{\varphi(i)}.$ (A.4)
Immediately from Definition A.4, one has $R_{\varphi}R_{\psi}=R_{\varphi\circ\psi}$, $R^{-1}_{\varphi}=R_{\varphi^{-1}}=R_{\varphi}^{\operatorname{T}}$, $\forall\varphi,\psi\in\text{Aut}(G)$. The following lemma outlines key properties for leveraging the symmetry of vertex-transitive graphs.
Lemma A.2 ([LS22, Lemma 4.1]).
  1. (i)
    $L$ commutes with $R_{\varphi}$ for every $\varphi\in\mathrm{Aut}(G)$.
  2. (ii)
    If $Y\in\mathbb{R}^{N\times N}$ commutes with $R_{\varphi}$ for every $\varphi\in\mathrm{Aut}(G)$, then $Y_{ii}=\frac{1}{N}\text{Tr}(Y)$, $\forall i\in[N]$.
  3. (iii)
    If $Y^{1},\cdots,Y^{N}\in\mathbb{R}^{N\times N}$ are such that $R_{\varphi}Y^{i}=Y^{\varphi(i)}R_{\varphi}$ holds for every $\varphi\in\mathrm{Aut}(G)$ and $i\in[N]$, then $Y_{ii}^{i}=Y_{jj}^{j}$, $\forall i,j\in[N]$.
A.1 Solving $F^{i}$ in terms of $P$
Our first step is to solve the Riccati equation (A.2) for $F_{t}^{i}$, treating $P$ as given, and represent the solution in terms of $P$. To this end, we consider the linear system for $(Y_{t}^{i},\Lambda_{t}^{i})$:
$\begin{bmatrix}\dot{Y}^{i}_{t}\\ \dot{\Lambda}^{i}_{t}\end{bmatrix}=\begin{bmatrix}P_{t}&-(\varepsilon-q^{2})Le_{i}e_{i}^{\operatorname{T}}L\\ e_{i}e_{i}^{\operatorname{T}}&-P_{t}\end{bmatrix}\begin{bmatrix}Y^{i}_{t}\\ \Lambda^{i}_{t}\end{bmatrix}$, (A.5)
with terminal conditions $Y^{i}_{T}=cLe_{i}e_{i}^{\operatorname{T}}L$, $\Lambda^{i}_{T}=I$. If the solution to equation (A.5) is found and $\Lambda^{i}_{t}$ is invertible for any $t\in[0,T]$, then $F^{i}_{t}=Y^{i}_{t}(\Lambda^{i}_{t})^{-1}$ is the solution to equation (A.2) [BJP00]. Under the condition $q^{2}=\varepsilon$, the equation for $Y_{t}^{i}$ becomes decoupled, resulting in $Y^{i}_{t}=\tilde{P}_{t}Y^{i}_{T}$, where
$\tilde{P}_{t}=e^{-\int_{t}^{T}P_{s}\,ds}.$ (A.6)
By substituting $Y^{i}_{t}=\tilde{P}_{t}Y^{i}_{T}$ into equation (A.5), we obtain an ODE for $\Lambda_{t}^{i}$: $\dot{\Lambda}^{i}_{t}=e_{i}e_{i}^{\operatorname{T}}\tilde{P}_{t}Y^{i}_{T}-P_{t}\Lambda^{i}_{t}$. The solution to this equation: $\Lambda^{i}_{t}=\tilde{P}^{-1}_{t}\left(\Lambda^{i}_{T}-\int_{t}^{T}\tilde{P}_{s}e_{i}e_{i}^{\operatorname{T}}\tilde{P}_{s}Y^{i}_{T}\,\mathrm{d}s\right)$. Therefore, incorporating the terminal conditions, the underlying linear system yields the following solution:
$\begin{cases}Y_{t}^{i}=c\tilde{P}_{t}Le_{i}e_{i}^{\operatorname{T}}L\\ \Lambda_{t}^{i}=\tilde{P}^{-1}_{t}\left(I-c\int_{t}^{T}\tilde{P}_{s}e_{i}e_{i}^{\operatorname{T}}\tilde{P}_{s}Le_{i}e_{i}^{\operatorname{T}}L\,\mathrm{d}s\right)\end{cases}.$ (A.7)
Assuming (A.2) and applying Lemma A.2(ii) gives,
$e_{i}^{\operatorname{T}}\tilde{P}_{s}Le_{i}=\frac{1}{N}\text{Tr}(\tilde{P}_{s}L).$ (A.8)
This simplifies $\Lambda_{t}^{i}$ to:
$\Lambda_{t}^{i}=\tilde{P}^{-1}_{t}\left(I-c\int_{t}^{T}\frac{\text{Tr}(\tilde{P}_{s}L)}{N}\tilde{P}_{s}e_{i}e_{i}^{\operatorname{T}}L\,\mathrm{d}s\right).$ (A.9)
Under assumption (A.3), the solution to equation (A.2) is provided as
$F^{i}_{t}=c\tilde{P}_{t}Le_{i}e_{i}^{\operatorname{T}}L\left(I-c\int_{t}^{T}\frac{\text{Tr}(\tilde{P}_{s}L)}{N}\tilde{P}_{s}e_{i}e_{i}^{\operatorname{T}}L\,\mathrm{d}s\right)^{-1}\tilde{P}_{t}.$ (A.10)
Up until now, we’ve derived a solution $F^{i}_{t}$ expressed in terms of $P$.
Remark A.3.
The condition $q^{2}=\varepsilon$ ensures that the underlying linear system for $(Y^{i}_{t},\Lambda^{i}_{t})$ is decoupled. Without this condition, we need to solve the full system (A.5), for which the existence of a closed-form solution remains unclear since, when $t\neq s$:
$\begin{bmatrix}P_{t}&-(\varepsilon-q^{2})Le_{i}e_{i}^{\operatorname{T}}L\\ e_{i}e_{i}^{\operatorname{T}}&-P_{t}\end{bmatrix}\text{ does not commute with }\begin{bmatrix}P_{s}&-(\varepsilon-q^{2})Le_{i}e_{i}^{\operatorname{T}}L\\ e_{i}e_{i}^{\operatorname{T}}&-P_{s}\end{bmatrix}$.
A.2 Expressing $P$ in terms of $\eta$
To further simplify equation (A.10), we define
$\Sigma^{i}_{t}:=\int_{t}^{T}\frac{\text{Tr}(\tilde{P}_{s}L)}{N}\tilde{P}_{s}e_{i}e_{i}^{\operatorname{T}}L\,\mathrm{d}s,$ (A.11)
and
$\eta^{i}_{t}:=e_{i}^{\operatorname{T}}L\left(I-c\Sigma^{i}_{t}\right)^{-1}\tilde{P}_{t}e_{i}.$ (A.12)
Thus $F_{t}^{i}$ admits the representation $F^{i}_{t}=c\tilde{P}_{t}Le_{i}e_{i}^{\operatorname{T}}L\left(I-c\Sigma^{i}_{t}\right)^{-1}\tilde{P}_{t}$ and satisfies $F^{i}_{t}e_{i}=c\eta^{i}_{t}\tilde{P}_{t}Le_{i}$. Under assumption (A.2), using Lemma A.2(i) and equation (A.11) leads to $R_{\varphi}\Sigma^{i}_{t}=\Sigma^{\varphi(i)}_{t}R_{\varphi}$, $\forall\varphi\in\mathrm{Aut}(G)$. With the additional assumption (A.4), we find:
$R_{\varphi}L\left(I-c\Sigma^{i}_{t}\right)^{-1}=L\left(I-c\Sigma^{\varphi(i)}_{t}\right)^{-1}R_{\varphi},\ \forall\varphi\in\mathrm{Aut}(G).$ (A.13)
By Lemma A.2(iii), $\eta^{i}_{t}$ becomes independent of the player index $i$. From this point, we will use the notation $\eta_{t}$ instead of $\eta_{t}^{i}$, as this quantity is shared by all players.
We now proceed to derive an expression for $P_{t}$ in terms of $\eta_{t}$. Substituting the relation $F^{i}_{t}e_{i}=c\eta_{t}\tilde{P}_{t}Le_{i}$ into equation (A.1) yields
$P_{t}=(a+q)L+c\eta_{t}\tilde{P}_{t}L,$ (A.14)
where $\tilde{P}$ is defined as in equation (A.6). It remains to solve the above equation for $P_{t}$.
Under assumption (A.1), any two symmetric matrices in the collection $\{L\}\cup\{P_{t},\tilde{P}_{t}:t\in[0,T]\}$ commute, which implies that this collection of matrices can be diagonalized simultaneously. Therefore, one can work with the spectra of the matrices rather than directly with the matrices themselves, providing greater convenience for calculations. Let $\rho^{1}_{t},\ldots,\rho^{N}_{t}$ be the eigenvalues of $P_{t}$ and $\lambda_{1},\ldots,\lambda_{N}$ be the eigenvalues of $L$, equation (A.14) becomes
$\rho^{k}_{t}=(a+q)\lambda_{k}+c\eta_{t}\lambda_{k}e^{-\int_{t}^{T}\rho^{k}_{s}\,\mathrm{d}s},\ \forall k\in[N].$ (A.15)
$e^{\int_{t}^{T}\rho^{k}_{s}-(a+q)\lambda_{k}\,\mathrm{d}s}=1+c\lambda_{k}\int_{t}^{T}\eta_{s}e^{-(T-s)(a+q)\lambda_{k}}\,\mathrm{d}s.$ (A.16)
$\rho^{k}_{t}=(a+q)\lambda_{k}+\frac{c\lambda_{k}\eta_{t}e^{-(T-t)(a+q)\lambda_{k}}}{1+c\lambda_{k}\int_{t}^{T}\eta_{s}e^{-(T-s)(a+q)\lambda_{k}}\,\mathrm{d}s}.$ (A.17)
$P_{t}=(a+q)L+c\eta_{t}Le^{-(T-t)(a+q)L}\left(I+\int_{t}^{T}c\eta_{s}e^{-(T-s)(a+q)L}\,\mathrm{d}s\;L\right)^{-1}.$ (A.18)
The above expression for $P_{t}$ in terms of $\eta_{t}$ is not entirely satisfactory because $\eta$, as defined in (A.12), depends on $\tilde{P}_{t}$, and therefore on $P_{t}$. Consequently, equation (A.18) does not provide an explicit construction for $P_{t}$. The next step (in Section A.3) is to find a way to characterize and construct $\eta$ based solely on model parameters. Once this is achieved, equation (A.18) can be used to construct $P_{t}$, as detailed in Section A.4.
A.3 The governing equation for $\eta$
According to the definition of $\eta$ in (A.12), the primary challenge in characterizing $\eta_{t}$ arises from the inverse term $\left(I-c\Sigma^{i}_{t}\right)^{-1}$. Combining equations (A.8) and (A.11) gives
$(\Sigma^{i}_{t})^{2}=\left[\int_{t}^{T}\left(\frac{\text{Tr}(\tilde{P}_{s}L)}{N}\right)^{2}\,\mathrm{d}s\right]\Sigma^{i}_{t}.$ (A.19)
$\left(I-c\Sigma^{i}_{t}\right)^{-1}=I+c\left[1-c\int_{t}^{T}\left(\frac{\text{Tr}(\tilde{P}_{s}L)}{N}\right)^{2}\,\mathrm{d}s\right]^{-1}\Sigma^{i}_{t}.$ (A.20)
Plugging it back into equation (A.12) and combining it with equation (A.8) yields
$\eta_{t}=\frac{g_{t}}{1-c\int_{t}^{T}g^{2}_{s}\,\mathrm{d}s},$ (A.21)
where $g_{t}$ is a scalar function defined as:
$g_{t}:=\frac{\text{Tr}(\tilde{P}_{t}L)}{N}.$ (A.22)
A property of $g_{t}$ will be repeatedly referred to later, so we derive it here. Direct calculations based on equation (A.14) gives:
$\text{Tr}(\tilde{P}_{t}L)=\frac{1}{c\eta_{t}}\left[\text{Tr}(P_{t})-(a+q)\text{Tr}(L)\right].$ (A.23)
Given $\text{Tr}(P_{t})=\sum_{k=1}^{N}\rho^{k}_{t}$, combining it with equation (A.17) yields:
$Ng_{t}=\text{Tr}(\tilde{P}_{t}L)=\sum_{k=1}^{N}\frac{\lambda_{k}e^{-(T-t)(a+q)\lambda_{k}}}{1+c\lambda_{k}\int_{t}^{T}\eta_{s}e^{-(T-s)(a+q)\lambda_{k}}\,\mathrm{d}s}.$ (A.24)
Back to equation (A.21), by multiplying both sides by $cg_{t}$, integrating in $t$, and taking the exponential, we obtain:
$e^{c\int_{t}^{T}g_{s}\eta_{s}\,\mathrm{d}s}=e^{\int_{t}^{T}\frac{cg^{2}_{s}}{1-c\int_{s}^{T}g^{2}_{u}\,\mathrm{d}u}\,\mathrm{d}s}=\frac{1}{1-c\int_{t}^{T}g^{2}_{s}\,\mathrm{d}s}.$ (A.25)
Combining equations (A.22) and (A.24), one has
$c\int_{t}^{T}g_{s}\eta_{s}\,\mathrm{d}s=\frac{1}{N}\sum_{k=1}^{N}\int_{t}^{T}\frac{c\lambda_{k}\eta_{s}e^{-(T-s)(a+q)\lambda_{k}}}{1+c\lambda_{k}\int_{s}^{T}\eta_{u}e^{-(T-u)(a+q)\lambda_{k}}\,\mathrm{d}u}\,\mathrm{d}s.$ (A.26)
Combining equations (A.25)–(A.26) and using similar arguments as done for equation (A.16) implies
$\frac{1}{1-c\int_{t}^{T}g^{2}_{s}\,\mathrm{d}s}=\left[\prod_{k=1}^{N}\left(1+c\lambda_{k}\int_{t}^{T}\eta_{s}e^{-(T-s)(a+q)\lambda_{k}}\,\mathrm{d}s\right)\right]^{\frac{1}{N}}.$ (A.27)
Plugging equations (A.22), (A.24) and (A.27) into equation (A.21) yields
$\eta_{t}=\frac{1}{N}\left(\sum_{k=1}^{N}\frac{\lambda_{k}e^{-(T-t)(a+q)\lambda_{k}}}{1+c\lambda_{k}\int_{t}^{T}\eta_{s}e^{-(T-s)(a+q)\lambda_{k}}\,\mathrm{d}s}\right)\\ \left[\prod_{k=1}^{N}\left(1+c\lambda_{k}\int_{t}^{T}\eta_{s}e^{-(T-s)(a+q)\lambda_{k}}\,\mathrm{d}s\right)\right]^{\frac{1}{N}}.$ (A.28)
Finally, rewriting equation (A.28) in the matrix form we obtain:
$\eta_{t}=\frac{1}{N}\text{Tr}\left[Le^{-(T-t)(a+q)L}\left(I+cL\int_{t}^{T}\eta_{s}e^{-(T-s)(a+q)L}\,\mathrm{d}s\right)^{-1}\right]\\ \left[\det\left(I+cL\int_{t}^{T}\eta_{s}e^{-(T-s)(a+q)L}\,\mathrm{d}s\right)\right]^{\frac{1}{N}}.$ (A.29)
Observe that equation (A.29) involves only the model parameters and the graph Laplacian $L$ without relying on $P_{t}$, which aligns with our goal of characterizing $\eta_{t}$. Nevertheless, equation (A.29) remains complex and does not easily allow for proving existence and uniqueness. To address this, we apply matrix calculus techniques. The lemma below plays an important role in formulating the governing equation for $\eta_{t}$.
Lemma A.4 ([PP08, Equation (41)]).
$\frac{\mathrm{d}\det A(t)}{\mathrm{d}t}=\text{Tr}\left(\mathrm{adj}[A(t)]\frac{\mathrm{d}A(t)}{\mathrm{d}t}\right),$ (A.30)
where $\mathrm{adj}[B]:=B^{-1}\det B$ is the adjugate matrix, well-defined for any nonsingular matrix $B$.
$\frac{\mathrm{d}\det A(X)}{\mathrm{d}X}=\frac{\text{Tr}\left(\mathrm{adj}[A(X)]\mathrm{d}A(X)\right)}{\mathrm{d}X}.$ (A.31)
The next corollary follows easily from Lemma A.31 and the chain rule of matrix calculus.
Corollary A.5.
$\frac{\mathrm{d}[\det(I+cLR(t))]^{\frac{1}{N}}}{\mathrm{d}t}=\frac{1}{N}\text{Tr}\left([I+cLR(t)]^{-1}cLR^{\prime}(t)\right)[\det(I+cLR(t))]^{\frac{1}{N}}.$ (A.32)
$\frac{\mathrm{d}[\det(I+cXL)]^{\frac{1}{N}}}{\mathrm{d}X}=\frac{1}{N}[\det(I+cXL)]^{\frac{1}{N}}(I+cXL)^{-1}cL.$ (A.33)
$R(t):=\int_{T-t}^{T}\eta_{s}e^{-(T-s)(a+q)L}\,\mathrm{d}s,$ (A.34)
then, based on equation (A.29) and Corollary A.5 and the definition of $Q$, $R(t)$ solves the following matrix-valued ODE:
$R^{\prime}(t)=\frac{1}{c}\text{Tr}\left[Q^{\prime}(R(t))e^{-t(a+q)L}\right]e^{-t(a+q)L},\quad R(0)=0,$ (A.35)
$R^{\prime}(T-t)=\eta_{t}e^{-(T-t)(a+q)L}$ (A.36)
provides the corresponding value of $\eta_{t}$.
Remark A.6 (Assumptions for heuristic derivations).
The following assumptions are made in the heuristic construction:
  1. (A.1)
  2. (A.2)
  3. (A.3)
    $\Lambda^{i}_{t}$ is invertible, $\forall t\in[0,T]$, $\forall i\in[N]$.
  4. (A.4)
A.4 The construction of the Markovian NE
Once $\eta_{t}$ has been constructed based on equations (4.3)–(4.4), the Markovian NE can be determined as follows. From equations (A.18) and (A.36), $P_{t}$ can be represented in terms of $R$ as follows:
$P_{t}=(a+q)L+R^{\prime}(T-t)\;cL[I+R(T-t)\;cL]^{-1}.$ (A.37)
which coincides with equation (4.6). By combining equations (A.10)–(A.11) and (A.20), we obtain
$F^{i}_{t}=\frac{c}{1-c\int_{t}^{T}\left(\frac{\text{Tr}(\tilde{P}(s)L)}{N}\right)^{2}\,\mathrm{d}s}\tilde{P}_{t}Le_{i}e_{i}^{\operatorname{T}}L\tilde{P}_{t}.$ (A.38)
The denominator of $F^{i}_{t}$ can be simplified using equations (A.22)–(A.23):
$1-c\int_{t}^{T}\left(\frac{\text{Tr}(\tilde{P}_{s}L)}{N}\right)^{2}\,\mathrm{d}s=\frac{1}{cN\eta^{2}_{t}}[\text{Tr}(P_{t})-(a+q)\text{Tr}(L)].$ (A.39)
Plugging equations (A.14) and (A.39) into equation (A.38) leads to the result (4.5).